Booster: Summary of PS ops: Bill,  Typical running – very little downtime, collimator study completed last Thursday (analyzing data), no planned shutdowns this week and no systems with pending issues/downtime.  Shielding assessment work progressing – when complete will allow for higher flux to BNB.

BNB: Tom, Fan continuing to die, think it is solid state (radiation), when it’s windy, radiation gets stuck in the air in MI12. Can try ramping BNB down if rates get too high. NuMI 6+6 studies affect BNB rate but not intensity.

Planned down time on April 6.

ANNIE: Installing electronics. Try readout electronics this week. Water next week.

MicroBooNE: GPS time firmware needs to be updated.

MiniBooNE: Fine.

Tom: Want a target scan. Can coordinate some timing adjustments with MIBS and A9.